OpenStudy (anonymous):

Convert the complex number 3+3i into its polar representation.

OpenStudy (anonymous):

Answer choices.

OpenStudy (anonymous):

draw a picture and you will see right away that the angle is \(\frac{\pi}{4}\) only thing left you need it \(r=\sqrt{3^2+3^2}=3\sqrt{2}\)

OpenStudy (anonymous):

oops i didn't look at your choices. apparently you are working with degrees, so it will be 45 degrees pick answer 2

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

yw